# ROSE: Register-assisted General Time Series Forecasting with Decomposed Frequency Learning

### v2

@@ Abstract
With the increasing collection of time series data from various domains, there arises a strong demand for general time series forecasting models pre-trained on a large number of time-series datasets to support a variety of downstream prediction tasks. Enabling general time series forecasting faces two challenges: how to obtain unified representations from multi-domian time series data, and how to capture domain-specific features from time series data across various domains for adaptive transfer in downstream tasks. To address these challenges, we propose a Register Assisted General Time Series Forecasting Model with Decomposed Frequency Learning (ROSE), a novel pre-trained model for time series forecasting. ROSE employs Decomposed Frequency Learning for the pre-training task, which decomposes coupled semantic information in time series with frequency-based masking and reconstruction to obtain unified representations across domains. We also equip ROSE with a Time Series Register, which learns to generate a register to capture domain-specific representations during pre-training and enhances domain-adaptive transfer by selecting related register tokens on downstream tasks. After pre-training on large-scale time series data, ROSE achieves state-of-the-art forecasting performance on 7 real-world benchmarks. Remarkably, it demonstrates competitive or superior few-shot and zero-shot abilities.
@@ 1 Introduction
Time Series Forecasting plays a pivotal role across numerous domains, such as energy, smart transportation, weather, and economics (Qiu et al., 2024). However, training specific models in deep learning for each dataset is costly and requires tailored parameter tuning, whose prediction accuracy may be limited due to data scarcity (Liu et al., 2024). One solution is to pre-train a general model on diverse time series datasets and fine-tune it with a few data for different downstream scenarios or direct predict without fine-tuning. Following this idea, foundation models for time series forecasting have recently raised growing attention, continually scaling up the pre-training datasets and model sizes to improve the generalization performance (Woo et al., 2024; Goswami et al., 2024; Ansari et al., 2024). However, an excessively large scale also causes increasing costs in training and inference, which may go against the original intention of a general model, especially in resource-constrained situations. In addition to scaling up, general time series forecasting models can also be designed from the perspectives of pre-training tasks and downstream transfer adaptation. From these two angles, we identify the following challenges.
Obtaining a unified representation from time series data across various domains is challenging. Time series from each domain involve complex temporal patterns, composed of multiple frequency components combined with each other (Zhou et al., 2022), which is frequency superposition. As shown in Figure 1(a), different frequency components contain distinct semantic information. For example, low and high-frequency components represent long-term trends and rapid variations, respectively (Zhang et al., 2022). Furthermore, different datasets exhibit diverse frequency distributions, and the significance of low-frequency and high-frequency components for time series modeling varies across domains(Zhang et al., 2024). As a result, large-scale time series data from different domains introduce even more complex temporal patterns and frequency diversity. Existing pre-training frameworks (Dong et al., 2024; Nie et al., 2022; Lee et al., 2023), such as masked modeling and contrastive learning, were proposed to learn a unified representation from time domain. However, these methods overlook the frequency diversity and complexity exhibited in heterogeneous time series that come from various domains, making it difficult to capture intricate patterns, thus limiting their generalization capabilities.
Adaptive transferring information from multi-domain time series to specific downstream scenarios presents a challenge. Multi-source time series data originate from various domains (Woo et al., 2024), whose data exhibit domain-specific information (Liu et al., 2024). Information from the same or similar domain as the target domain is useful for improving the model’s effectiveness in the target task (Chen et al., 2023a). However, as shown in Figure 1(a), existing time series pre-training frameworks (Woo et al., 2024; Liu et al., 2024; Zhou et al., 2024) focus mainly on learning generalized representation during pre-training and overlook domain-specific representation. Thus, they only transfer the same generalized representation to different target domains, called direct transfer, which limits the model’s effectiveness in specific downstream tasks. Therefore, it is necessary to learn domain-specific information during pre-training and adaptively transfer the specific representations to each target domain, called adaptive transfer. Realizing adaptive transfer poses two difficulties: 1) capturing domain-specific information in pre-training. 2) adaptive use of domain-specific information in various downstream tasks.
@@
Figure 1: (a) Pre-training on multi-domain datasets that exhibit combined frequency. Existing general time series forecasting models only extract generalized representations for direct transfer to various downstream target domains. We propose to learn generalized and specific representations during pre-training, and adaptively transfer them to each target domain. (b) The t-SNE visualization of the hidden representations after direct transfer and adaptive transfer: In direct transfer, representations of different domains are mixed, but in adaptive transfer, they show a clear clustering pattern.
@@
To address these challenges, we propose a register assisted general time series forecasting model with decomposed frequency learning (ROSE). First, we propose Decomposed Frequency Learning that learns generalized representations to solve the issue with coupled semantic information. We decompose individual time series using the Fourier transform with a novel frequency-based masking method, and then convert it back to the time domain to obtain decoupled time series for reconstruction. It makes complex temporal patterns disentangled, thus benefiting the model to learn generalized representations. Second, we introduce Time Series Register (TS-Register) to learn domain-specific information in multi-domain data. By setting up a register, we generate register tokens to learn each domain-specific information during pre-training. In a downstream scenario, the model adaptively selects Top-K vectors from the register that are close to the target domain of interest. During fine-tuning, we adjust the selected register tokens with a novel learnable low-rank matrix, which complements target-specific information to perform more flexible adaptive transfer. As shown in Figure 1(b), adaptive transfer successfully utilizes domain-specific information in multi-domain time series, which contributes to the model’s performance in target tasks. The contributions are summarized as follows:
@@ 2.1 Traditional time series forecasting
The statistical time series forecasting models like ARIMA (Box and Jenkins, 1968), despite their theoretical support, are limited in modeling nonlinearity. With the rise of deep learning, many RNN-based models (Cirstea et al., 2019; Wen et al., 2017; Salinas et al., 2020) have been proposed, modeling the sequential data with an autoregressive process. CNN-based models (Luo and Wang, 2024; Liu et al., 2022a) have also received widespread attention due to their ability to capture local features. MICN (Wang et al., 2022) utilizes TCN to capture both local and global features, while TimesNet (Wu et al., 2022) focuses on modeling 2D temporal variations. However, both RNNs and CNNs struggle to capture long-term dependencies. Transformer-based models (Zhou et al., 2022; Nie et al., 2022; Wu et al., 2021; Liu et al., 2023; Chen et al., 2024), with their attention mechanism, can capture long dependencies and extract global information, leading to widespread applications in long-time series prediction. However, this case-by-case paradigm requires meticulous hyperparameter design for different datasets, and its predictive performance can also be affected by data scarcity.
2.2 Pre-trained models for times seires forecasting
LLMs for time series forecasting: Recent studies have shown that Large Language Models (LLMs) can enhance time series forecasting with limited fine-tuning (Zhou et al., 2024), prompting (Jin et al., 2024; Cao et al., 2024), and modality alignment (Jin et al., 2024). GPT4TS (Zhou et al., 2024) fine-tunes a subset of LLM parameters, and achieves competitive performance by leveraging pre-trained text knowledge. TimeLLM (Jin et al., 2024) transforms time series into text to align with LLM representations. Despite their positive impact, LLMs’ high inference costs limit their practicality.
Time series foundation model: Pre-training with multiple sources time series has recently received widespread attention (Rasul et al., 2023; Dooley et al., 2024; Garza and Mergenthaler-Canseco, 2023; Kamarthi and Prakash, 2023). MOMENT (Goswami et al., 2024) and MOIRAI (Woo et al., 2024) adopt a BERT-style pre-training approach, while Timer (Liu et al., 2024), Chronos (Ansari et al., 2024) and TimsFM (Das et al., 2023a) use a GPT-style pre-training approach, giving rise to improved performance in time series prediction. However, the above methods overlook domain-specific information from multi-source data, thus limiting the performance of the models. Different from previous approaches, ROSE pre-trains on large-scale data from various domains and it considers both generalized representations and domain-specific information, which facilitates flexible adaptive transfer in downstream tasks.
@@ 3 Methodology
Problem Definition. Given a multivariate time series $\mathbf{X}_{t}=\{\mathbf{x}_{t-L:t}^{i}\}_{i=1}^{C}$, where each $\mathbf{x}_{t-L:t}^{i}\in\mathbb{R}^{L}$ is a sequence of observations. $L$ denotes the look-back window and $C$ denotes the number of channels. The forecasting task is to predict the future values $\mathbf{\hat{Y}}_{t}=\{\mathbf{\hat{x}}_{t:t+F}^{i}\}_{i=1}^{C}$, where $F$ denotes the forecast horizon. $\mathbf{Y}_{t}=\{\mathbf{x}_{t:t+F}^{i}\}_{i=1}^{C}$ is the ground truth of future.
The general time series forecasting model is pre-trained with multi-source datasets $\mathbf{D}_{\text{pre-train}}=\{(\mathbf{X}_{t}^{j},\mathbf{Y}_{t}^{j})\}^{N}_{j=1}$, where $N$ is the number of datasets. For the downstream task, the model is fine-tuned with a training dataset $\mathbf{D}_{\text{train}}=\{(\mathbf{X}_{t}^{\text{train}},\mathbf{Y}_{t}^{\text{train}})\}$, and is tested with $\mathbf{D}_{\text{test}}=\{(\mathbf{X}_{t}^{\text{test}},\mathbf{Y}_{t}^{\text{test}})\}$ to predict $\mathbf{\hat{Y}}_{t}^{\text{test}}$, where $\mathbf{D}_{\text{pre-train}}$, $\mathbf{D}_{\text{train}}$ and $\mathbf{D}_{\text{test}}$ are pairwise disjoint. Alternatively, the model could be directly tested using $\mathbf{D}_{\text{test}}$ without fine-tuning with $\mathbf{D}_{\text{train}}$ to predict $\mathbf{\hat{Y}}_{t}^{\text{test}}$.
@@ 3.1 Architecture
Figure 2: The model architecture of ROSE.
The time series forecasting paradigm of ROSE contains two steps: pre-training and fine-tuning. First, ROSE is pre-trained on large-scale datasets from various domains, with two pre-training tasks: reconstruction and prediction. We set up the reconstruction task to help the model understand time series comprehensively and set the prediction task to enhance the model’s few-shot and zero-shot abilities. Then, ROSE is fine-tuned with a target dataset in the downstream scenario.
As shown in Figure 2, ROSE employs the encoder-decoder architecture to model time series. The backbone consists of multiple Transformer layers to process sequential information and effectively capture temporal dependencies (Vaswani et al., 2017). The reconstruction decoder and prediction decoder use the same structure as the Transformer encoder. They are used for reconstruction and prediction tasks respectively. ROSE is pre-trained in a channel-independent way, which is widely used in time series forecasting (Nie et al., 2022).
@@ 3.2 Decomposed frequency learning
Multi-frequency masking. As shown in the green part of Figure 3, given a time series $\mathbf{x}\in\mathbb{R}^{L}$, we utilize the Real Fast Fourier Transform (rFFT) (Brigham and Morrow, 1967) to transform it into the frequency domain, giving rise to $\mathbf{x}_{\mathrm{freq}}\in\mathbb{C}^{L/2+1}$.
@@
Representation learning. As shown in the yellow part of Figure 3, after obtaining the $K_{\mathrm{f}}$ masked sequences $\mathbf{X}_{\mathrm{mask}}$, we divide each sequence $\mathbf{x}_{\mathrm{mask}}^{i}$ into $P$ non-overlapping patches, and use a linear layer to transforming them into $P$ patch tokens, and thus we get $\mathcal{X}_{\mathrm{mp}}=\{\mathbf{X}_{\mathrm{mp}}^{i}\}_{i=1}^{K_{\mathrm{f}}}$ to capture general information, where each $\mathbf{X}_{\mathrm{mp}}^{i}\in\mathbb{R}^{P\times D}$, and $D$ is the dimension for each patch token. We replicate the register tokens $\mathbf{X}_{\mathrm{u}}$ $K_{\mathrm{f}}$ times to get $\mathcal{X}_{\mathrm{u}}\in\mathbb{R}^{K_{\mathrm{f}}\times N_{\mathrm{r}}\times D}$, where $\mathbf{X}_{\mathrm{u}}\in\mathbb{R}^{N_{\mathrm{r}}\times D}$ is obtained by inputting the original sequence into the TS-Register, as detailed in Section 3.3. Then, we concatenate the patch tokens $\mathcal{X}_{\mathrm{mp}}$ with the register tokens $\mathcal{X}_{\mathrm{u}}$, and feed them into the Transformer encoder to obtain the representation of each masked series. These representations are then aggregated to yield a unified representation $\mathbf{S}_{\mathrm{m}}\in\mathbb{R}^{(N_{\mathrm{r}}+P)\times D}$. The aggregator is the averaging operation.
$\mathbf{S}_{\mathrm{m}}=\mathrm{Aggregator(Encoder(Concatenate(\mathcal{X}_{mp},\mathcal{X}_{u})))}.$ (4)
@@
$\mathcal{L}_{\mathrm{reconstruction}}=\mathrm{||\mathbf{x}-\hat{\mathbf{x}}||_{2}^{2}}.$ (5)
@@ 3.3 Time series register
By decomposed frequency learning, we can obtain the general representations, and additionally, we propose the TS-Register that learns register tokens as the domain-specific information from the multi-domain datasets for adaptive transfer. It clusters domain-specific information from the multi-domain datasets into register tokens and stores such domain-specific information in the register during pre-training. Then, it adaptively selects domain-specific information from the register via a Top-K selection strategy to enhance the performance in the target domain. A novel learnable low-rank matrix is proposed to set to complement the downstream dataset-specific information through fine-tuning.
@@
To update the cluster center vectors in the register $\mathbf{E}$ that represents the domain information of the pre-trained datasets, we set the loss function shown in Equation 6 that minimizes the distance between the data-dependent embedding $\mathbf{x}_{\mathrm{e}}$ and the cluster center $\mathbf{e}_{\delta}$. To solve the problem that the gradient of the $\arg\operatorname{min}$ function cannot be backpropagated, we use the stop gradient operation to pass the gradient of $\mathbf{e}_{\delta}$ directly to $\mathbf{x}_{\mathrm{e}}$.
@@ 3.4 Training
To improve the model’s prediction ability in zero-shot and few-shot setting, we co-train supervised prediction with self-supervised reconstruction that uses multi-frequency masking to learn unified features that are more applicable to the downstream prediction task. We normalize time series by employing the REVIN (Kim et al., 2021) that is commonly used by the state-of-the-art time series models, which first normalizes each input sample and subsequently applies inverse normalization to recover the model’s output.
Prediction task. The input time series $\mathbf{x}\in\mathbb{R}^{L}$ is sliced into $P$ non-overlapping patches and then mapped to $\mathbf{X}_{\mathrm{p}}\in\mathbb{R}^{P\times D}$. Based on common forecasted needs (Qiu et al., 2024), we set up four prediction heads mapping to prediction lengths of $\{96,192,336,720\}$ to accomplish the prediction task. Patch tokens $\mathbf{X}_{\mathrm{p}}$ are concatenated with the register tokens $\mathbf{X}_{\mathrm{u}}$ and then successively fed into the Transformer encoder to yield the representation $\mathbf{S}\in\mathbb{R}^{(N_{\mathrm{r}}+P)\times D}$:

### v3

# Towards a General Time Series Forecasting Model with Unified Representation and Adaptive Transfer

@@ Abstract
With the growing availability of multi-domain time series data, there is an increasing demand for general forecasting models pre-trained on multi-source datasets to support diverse downstream prediction scenarios. Existing time series foundation models primarily focus on scaling up pre-training datasets and model sizes to enhance generalization performance. In this paper, we take a different approach by addressing two critical aspects of general forecasting models: (1) how to derive unified representations from heterogeneous multi-domain time series data, and (2) how to effectively capture domain-specific features to enable adaptive transfer across various downstream scenarios. To address the first aspect, we propose Decomposed Frequency Learning as the pre-training task, which leverages frequency-based masking and reconstruction to decompose coupled semantic information in time series, resulting in unified representations across domains. For the second aspect, we introduce the Time Series Register, which captures domain-specific representations during pre-training and enhances adaptive transferability to downstream tasks. Our model achieves the state-of-the-art forecasting performance on seven real-world benchmarks, demonstrating remarkable few-shot and zero-shot capabilities.
Machine Learning, ICML
@@ 1 Introduction
Time series forecasting plays a crucial role in various domains, including energy, smart transportation, weather, and economics (Qiu et al., 2024; Wu et al., 2025a; Qiu et al., 2025b). However, training deep learning models for each specific dataset is resource-intensive and requires tailored parameter tuning. This approach often suffers from limited prediction accuracy due to data scarcity (Liu et al., 2024; Wu et al., 2021b, 2024, 2025c). A promising solution is to pre-train a general model on diverse time series datasets, which can then be fine-tuned with minimal data for different downstream scenarios or even used directly without fine-tuning. Following this idea, foundation models for time series forecasting have gained significant attention. Recent efforts have focused on scaling up pre-training datasets and model sizes to enhance generalization performance (Woo et al., 2024; Goswami et al., 2024; Ansari et al., 2024). However, excessive scaling introduces high computational costs during training and inference, undermining the practicality of general models, particularly in resource-constrained settings. Beyond scaling, the design of general time series forecasting models can also be approached through pre-training tasks and downstream transfer adaptation. From these two perspectives, we identify the following challenges.
@@
Figure 1: (a) Pre-training on multi-domain datasets that exhibit combined frequency. Existing general time series forecasting models only extract generalized representations for direct transfer to various downstream target domains. We propose to learn generalized and specific representations during pre-training, and adaptively transfer them to each target domain. (b) The t-SNE visualization of the hidden representations after direct transfer and adaptive transfer: In direct transfer, representations of different domains are mixed, but in adaptive transfer, they show a clear clustering pattern. The detailed experiment setting is in the Appendix A.1.3
Obtaining a unified representation from time series data across various domains is challenging. Time series from each domain involve complex temporal patterns, composed of multiple frequency components combined with each other (Zhou et al., 2022; Wu et al., 2025b), which is frequency superposition. As shown in Figure 1(a), different frequency components contain distinct semantic information. For example, low and high-frequency components represent long-term trends and rapid variations, respectively (Zhang et al., 2022). Furthermore, different datasets exhibit diverse frequency distributions, and the significance of low-frequency and high-frequency components for time series modeling varies across domains(Zhang et al., 2024). As a result, large-scale time series data from different domains introduce even more complex temporal patterns and frequency diversity. Existing pre-training frameworks (Dong et al., 2024; Nie et al., 2022; Lee et al., 2023), such as masked modeling and contrastive learning, were proposed to learn a unified representation from time domain. However, these methods overlook the frequency diversity and complexity exhibited in heterogeneous time series that come from various domains, making it difficult to capture intricate patterns, thus limiting their generalization capabilities.
Adaptive transferring information from multi-domain time series to specific downstream scenarios presents a challenge. Multi-source time series data originate from various domains (Woo et al., 2024), whose data exhibit domain-specific information (Liu et al., 2024; Miao et al., 2024; Zhao et al., 2023; Guo et al., 2014). Information from the same or similar domain as the target domain is useful for improving the model’s effectiveness in the target task (Chen et al., 2023a). However, as shown in Figure 1(a), existing time series pre-training frameworks (Woo et al., 2024; Liu et al., 2024; Zhou et al., 2024) primarily focus on learning generalized time series representations during pre-training while overlooking domain-specific representations, called direct transfer. While generalized representations are essential, directly transferring them to specific downstream tasks without incorporating domain-specific information leaves room for improvement. Therefore, it is necessary to learn domain-specific information during pre-training and adaptively transfer the specific representations to target domain, called adaptive transfer. Realizing adaptive transfer poses two difficulties: 1) capturing domain-specific information in pre-training. 2) adaptive use of domain-specific information in various downstream tasks.
@@
To address these challenges, we propose a register assisted general time series forecasting model with decomposed frequency learning, namely ROSE. First, we propose Decomposed Frequency Learning that learns generalized representations to solve the issue with coupled semantic information. We decompose individual time series using the Fourier transform with a novel frequency-based masking method, and then convert it back to the time domain to obtain decoupled time series for reconstruction. It makes complex temporal patterns disentangled, thus benefiting the model to learn generalized representations. Second, we introduce Time Series Register (TS-Register) to learn domain-specific information in multi-domain data. By setting up a register, we generate register tokens to learn each domain-specific information during pre-training. In a downstream scenario, the model adaptively selects Top-K vectors from the register that are close to the target domain of interest. During fine-tuning, we adjust the selected register tokens with a novel learnable low-rank matrix, which complements target-specific information to perform more flexible adaptive transfer. As shown in Figure 1(b), adaptive transfer successfully utilizes domain-specific information in multi-domain time series, which contributes to the model’s performance in target tasks. The contributions are summarized as follows:
@@ 2.1 Traditional Time Series Forecasting
The statistical time series forecasting models like ARIMA (Box & Jenkins, 1968), despite their theoretical support, are limited in modeling nonlinearity. With the rise of deep learning, many RNN-based models (Cirstea et al., 2019; Wen et al., 2017; Salinas et al., 2020) have been proposed, modeling the sequential data with an autoregressive process. CNN-based models (Luo & Wang, 2024; Liu et al., 2022b) have also received widespread attention due to their ability to capture local features. MICN (Wang et al., 2022) utilizes TCN to capture both local and global features, while TimesNet (Wu et al., 2022) focuses on modeling 2D temporal variations. However, both RNNs and CNNs struggle to capture long-term dependencies. Transformer-based models (Zhou et al., 2022; Nie et al., 2022; Wu et al., 2021a; Liu et al., 2023; Chen et al., 2024), with their attention mechanism, can capture long dependencies and extract global information, leading to widespread applications in long-time series prediction. However, this case-by-case paradigm requires meticulous hyperparameter design for different datasets, and its predictive performance can also be affected by data scarcity.
2.2 Time Series Forecasting Foundation Model
Pre-training with multiple sources time series has recently received widespread attention (Rasul et al., 2023; Dooley et al., 2024; Garza & Mergenthaler-Canseco, 2023; Kamarthi & Prakash, 2023). MOMENT (Goswami et al., 2024) and MOIRAI (Woo et al., 2024) adopt a BERT-style pre-training approach, while Timer (Liu et al., 2024), Chronos (Ansari et al., 2024) and TimsFM (Das et al., 2023a) use a GPT-style pre-training approach, giving rise to improved performance in time series prediction. However, the above methods overlook domain-specific information from multi-source data, thus limiting the performance of the models. Different from previous approaches, ROSE pre-trains on large-scale data from various domains and it considers both generalized representations and domain-specific information, which facilitates flexible adaptive transfer in downstream tasks.
@@ 3.1 Architecture
As illustrated in Figure 2, ROSE adopts an encoder-decoder architecture for time series modeling. Its backbone comprises multiple Transformer layers, which effectively process sequential information and capture temporal dependencies (Vaswani et al., 2017). Both the reconstruction decoder and prediction decoder share the same structure as the Transformer encoder and are designed for reconstruction and prediction tasks, respectively. The reconstruction task enables the model to gain a comprehensive understanding of time series, while the prediction task enhances its few-shot and zero-shot capabilities. ROSE is pre-trained in a channel-independent way, which is widely used in time series forecasting (Nie et al., 2022).
@@ 3.2 Decomposed Frequency Learning
Multi-frequency masking. As shown in the green part of Figure 3, given a time series $\mathbf{x}\in\mathbb{R}^{L}$, we utilize the Real Fast Fourier Transform (rFFT) (Brigham & Morrow, 1967) to transform it into the frequency domain, giving rise to $\mathbf{x}_{\mathrm{freq}}\in\mathbb{C}^{L/2+1}$.
@@
Representation learning. As shown in the yellow part of Figure 3, after obtaining the $K_{\mathrm{f}}$ masked sequences $\mathbf{X}_{\mathrm{mask}}$, we divide each sequence $\mathbf{x}_{\mathrm{mask}}^{i}$ into $P$ non-overlapping patches, and use a linear layer to transforming them into $P$ patch tokens, and thus we get $\mathcal{X}_{\mathrm{mp}}=\{\mathbf{X}_{\mathrm{mp}}^{i}\}_{i=1}^{K_{\mathrm{f}}}$ to capture general information, where each $\mathbf{X}_{\mathrm{mp}}^{i}\in\mathbb{R}^{P\times D}$, and $D$ is the dimension for each patch token. We replicate the register tokens $\mathbf{X}_{\mathrm{u}}$ $K_{\mathrm{f}}$ times to get $\mathcal{X}_{\mathrm{u}}\in\mathbb{R}^{K_{\mathrm{f}}\times N_{\mathrm{r}}\times D}$, where $\mathbf{X}_{\mathrm{u}}$ is obtained by inputting the original sequence into the TS-Register, as detailed in Section 3.3. Then, we concatenate the patch tokens $\mathcal{X}_{\mathrm{mp}}$ with the register tokens $\mathcal{X}_{\mathrm{u}}$, and feed them into the Transformer encoder to obtain the representation of each masked series. These representations are then averaging aggregated to yield a unified representation $\mathbf{S}_{\mathrm{m}}\in\mathbb{R}^{(N_{\mathrm{r}}+P)\times D}$.
$\mathbf{S}_{\mathrm{m}}=\mathrm{Aggregator(Encoder(Concat(\mathcal{X}_{mp},\mathcal{X}_{u})))}.$ (4)
@@
[…]
@@ 3.3 Time Series Register
By decomposed frequency learning, we can obtain the general representations. Additionally, we propose the TS-Register that learns domain-specific information from the multi-domain datasets for adaptive transfer. It clusters domain-specific information from the multi-domain datasets into register tokens and stores such domain-specific information in the register during pre-training. Then, it adaptively selects domain-specific information from the register via a Top-K selection strategy to enhance the performance in the target domain. A novel learnable low-rank matrix is proposed to set to complement the downstream dataset-specific information through fine-tuning.
@@
To update the cluster center vectors in the register $\mathbf{E}$ that represents the domain information of the pre-trained datasets, we set the loss function shown in Equation 6 that minimizes the distance between the embedding $\mathbf{x}_{\mathrm{e}}$ and the cluster center $\mathbf{e}_{\delta}$. To solve the problem that the gradient of the $\arg\operatorname{min}$ function cannot be backpropagated, we use the stop gradient operation to pass the gradient of $\mathbf{e}_{\delta}$ directly to $\mathbf{x}_{\mathrm{e}}$.
@@ 3.4 Training
To improve the prediction performance in zero-shot and few-shot settings, we co-train supervised prediction with self-supervised reconstruction that uses multi-frequency masking to learn unified features that are more applicable to the downstream prediction task.
Prediction task. The input time series $\mathbf{x}\in\mathbb{R}^{L}$ is sliced into $P$ non-overlapping patches and then mapped to $\mathbf{X}_{\mathrm{p}}\in\mathbb{R}^{P\times D}$. Based on common forecasting needs (Qiu et al., 2024), we set up four prediction heads mapping to prediction lengths of $\{96,192,336,720\}$ to accomplish the prediction task. Patch tokens $\mathbf{X}_{\mathrm{p}}$ are concatenated with the register tokens $\mathbf{X}_{\mathrm{u}}$ and then successively fed into the Transformer encoder to yield the representation $\mathbf{S}\in\mathbb{R}^{(N_{\mathrm{r}}+P)\times D}$: